18. A 50/50 blend of engine coolant and water (by volume) is usually used in an automobile\'s engine cooling system. If your car\'s cooling system holds 6.20 gallons, what is the boiling point of the solution? Make the following assumptions in your calculation; at normal filling conditions, the densities of engine coolant and water are 1.11 g/mL and 0.998 g/mL respectively. Assume that the engine coolant is pure ethylene glycol (HOCH2CH2OH), which is non-ionizing and non-volatile, and that the pressure remains constant at 1.00 atm. Also, you\'ll need to look up the boiling-point elevation constant for water. ? _______ deg C
Answer #1
given capacity 6.2 gallons

3.1 coolant= 11.7335 lit
weight=12906.85 gm
moles=weight/62 =208.175
weight of water=3.1*3.875*0.998
molality= moles/weight of water in kg =17.3646
kb=0.512 c/m
boiling point =100+17.3646*0.512=108.89 c
Answer #3

3.1 coolant= 11.7335 lit
weight=12906.85 gm
moles=weight/62 =208.175
weight of water=3.1*3.875*0.998
molality= moles/weight of water in kg =17.3646
kb=0.512 c/m
boiling point =100+17.3646*0.512=108.89 c